OpenStudy (anonymous):

The subject is Linear Equations. Solve for x. (Please show your work.) 0.12(y - 6) + 0.06y = 0.08y - 0.7

OpenStudy (anonymous):

Errr, solve for y, my apologies.

OpenStudy (katrinakaif):

Your in college..? I expected your problems to be much more challenging O_o

OpenStudy (anonymous):

I'm in Math 65. I forgot essentially all there is to Algebra and its concepts.

OpenStudy (mimi_x3):

y=0.2 Is this college lvl!? gosh its easy

OpenStudy (katrinakaif):

Distrbute 0.12 to (y) and (-6) and add 0.06y with the Y variables only..Tell me what you get at the end.

OpenStudy (anonymous):

0.2. Thank you.

OpenStudy (katrinakaif):

So when distrbuted it would equal: 0.12y - 0.72 + 0.06y Collect like terms: 0.12y + 0.06y - 0.72 0.18y - 0.72 = 0.08y - 0.7 0.18y - 0.08y = -0.7 + 0.72 0.1y = 0.02 y = 0.02/0.1 y = 0.2